Calculate the work done when an object of mass, 50kg is moved through a distance of 5m. (g=10m/s).
1
Expert's answer
2020-04-02T10:50:23-0400

Work done ="F.d=Fd\cos\theta"

where F is the force and d is the distance moved and "\theta" is the angle between the force and the displacement vector

here mg is force acting vertically downwards and displacement vector is perpendicular to mg

hence cos 900 which will make the work done by the gravitational force =0